OpenStudy (anonymous):

Write an equation for a parabola in which the set of all points in the plane are equidistant from the focus and line. F(–2, 0); x = 2

OpenStudy (anonymous):

can anyone help im stuck on this last question

OpenStudy (anonymous):

these are my options below: x=1/8y^2 x=-1/8y^2 x=-1/4y^2 x=1/4y^2

OpenStudy (anonymous):

im thinking A

OpenStudy (anonymous):

or B